# Iterative Repair with Weak Verifiers for Few-shot Transfer in KBQA with Unanswerability

### v2

@@ Appendix A Appendix
A.1 KBQA Problem Formulations
We begin by defining the few-shot transfer learning setting for Knowledge Base Question Answering (KBQA) with unanswerability, and then present our solution for it.
A.1.1 KBQA: Detailed Background
A Knowledge Base $G$ consists of a schema with data stored under it. The schema consists of entity types $T$ and binary relations $R$ defined over pairs of types. The data consists of entities $E$ as instances of types $T$, and triples or facts $F\subseteq E\times R\times E$. We are given a target Knowledge Base $G^{t}$ (consisting of entity types $T^{t}$, relations $R^{t}$, entities $E^{t}$ and facts $F^{t}$) and a natural language question $q^{t}$, and the goal is to generate a structured query or a logical form $l^{t}$, which when executed over $G^{t}$ returns a non-empty answer $A^{t}$ for the question $q^{t}$.
A.1.2 Few Shot Transfer Learning for KBQA
In few-shot transfer learning, we are provided with target few-shots $D^{t}$ containing tens of labeled training examples of questions and logical forms in the target domain. In addition, we are given access to a related source domain. This has a source knowledge base $G^{s}$ consisting of its own types $T^{s}$, relations $R^{s}$, entities $E^{s}$ and facts $F^{s}$, and a much larger source training set $D^{s}$ containing thousands of labeled training examples of questions and corresponding logical forms. The source and target tasks may differ significantly. First, the data and schema of the knowledge bases $G^{t}$ and $G^{s}$ and the domains they cover may be different. Secondly, the distributions of questions and logical forms defined over the KBs may be different in $D^{t}$ and $D^{s}$.
A.1.3 KBQA with Unanswerability
A natural language question q is said to be answerable for a Knowledge Base $G^{t}$ if it has a corresponding logical form $l$ which when executed over $G^{t}$ returns a non-empty answer $A$. In contrast, a question $q$ is unanswerable for $G^{t}$, if it either (a) does not have a corresponding logical form that is valid for $G^{t}$, or (b) it has a valid logical form $l$ for $G^{t}$, but which on executing returns an empty answer. The first case indicates that $G^{t}$ is missing some schema element necessary for capturing the semantics for $q$. The second case indicates that the types $T^{s}$, relations $R^{s}$ is sufficient for $q$, but $G^{t}$ is missing some necessary data elements for answering it. In the KBQA with unanswerability task, given a question $q$, if it is answerable, the model needs to output the corresponding logical form $l$. and the non-empty answer $A$ entailed by it, and if it is unanswerable, the model either needs to output $NK$ (meaning No Knowledge) for the logical form, or a valid logical form $l$ with $NA$ (meaning No Answer) as the answer.
Different Types of Unanswerability:
@@ A.1.3 KBQA with Unanswerability
A.1.4 Few-shot Transfer Learning for KBQA with Unanswerability
This setting is a generalization of few shot transfer learning for KBQA, wherein we no longer assume that the test question is answerable with respect to the Knowledge Base $G^{t}$. The training set $D^{s}$ contains thousands of answerable questions along with their corresponding logical form. The target few shots $D^{t}$ contain both answerable as well as unanswerable questions, along with their corresponding logical forms, if they exist.
A.2 Error Checks
@@
(V3) Question Logical form Disagreement:
FUn performs equivalence check between $l$ and $q$ using a novel multi-stage LLM pipeline. (i) The variable names in $l$ are first naturalized to $l^{n}$ considering $q$ and preserving semantics, e.g. by replacing ’?x’ with ’?actor’. (ii) $l^{n}$ is back-translated into a natural language question $q^{b}$. (iii) $q^{b}$ is finally checked for semantic equivalence with $q$. The first two steps are performed using zero-shot prompting, while the last is performed using few-shots constructed using the target few-shots $D^{t}$. More details about few-shot construction are in Sec. A.8.2. The feedback mentions lack of equivalence as the type of error.
@@
A.3 EM-s: Automated check for equivalence of SPARQL programs
@@
As has been observed in Patidar et al. (2023), answer evaluation by itself is not a robust measure for evaluation of KBQA models when the dataset contains unanswerability. Traditional KBQA models that generate s-expressions can be evaluated using EM, which checks for logical form equivalence between two logical forms, since it is possible to compare equivalence between two s-expressions efficiently. However, FUn-FuSIC generates sparql queries instead. Directly comparing program equivalence between two sparql queries is an undecideable problem  [^6] https://users.dcc.uchile.cl/ cgutierr/papers/expPowSPARQL.pdf. Patidar et al. (2024) suggests a semi-automatic strategy for comparison of sparql queries. We propose a completely automatic metric for sparql equivalence check. Two sparql queries are equivalent by the EM-s check if (a) the relations occuring in the two queries are same. (b) the entities occuring in the two queries are the same (c) the answer set obtained by executing the queries over the KB are the same. Note that the EM-s check is necessary, but not sufficient for two sparql queries to be equivalent.
Since these are a necessary but not sufficient condition for logical form equivalence, we compared EM-s with EM, where both are applicable and found $>98\%$ agreement.
A.4 Performance across different categories of unanswerability
@@
We dive deeper into the different categories of unanswerability, as explained in Patidar et al. (2023). There exist 2 broad categories of unanswerability- schema level unanswerability(absence of knowledge in terms of KB ontology or entities required to construct the logical form) and data level unanswerability(absence of facts or intermediate entities of the logical form path on the KB).
@@
We expect that (a) due to the poor ability of supervised models to generalize in transfer learning settings, RetinaQA will be struggle to generate correct logical forms for data level unanswerable questions (b) due to the strong generalization ability of FuSIC-KBQA, it should be able to perform well for data level unanswerable questions. However, since it is biased to return incorrect logical forms instead of abstaining from returning a logical form, it will perform poorly at identifying schema level unanswerable questions. (c) FUn-FuSIC should be able to maintain the performance of FuSIC-KBQA on data level unanswerable questions to a large extent, while significantly improving the performance on schema level unanswerable questions.
Based on performance on the WebQSP $\rightarrow$ GrailQAbility and WebQSP $\rightarrow$ GraphQAbility datasets, we validate the that the trends are indeed as expected.
A.5 Cost Analysis
@@
FuSIC-KBQA, as well as the adapted versions of FuSIC-KBQA, such as U-FuSIC and FUn-FuSIC rerank the classes, relations and paths. The total cost for reranking for one question is \$0.16.
The cost for generation of logical form from a prompt with 5 in-context examples is \$0.16.
Thus, the approximate cost for inference of one question by U-FuSIC is \$0.32.
@@
The cost for generation of logical form from a prompt with 0 in-context examples is \$0.04. The cost of checking whether two natural language questions are equivalent or not, using few-shot exemplars and chain of thought prompting is also \$0.04.
The approximate cost of inference of one question by FUn-FuSIC varies between \$0.24 and \$0.48. The average cost over 50 randomly sampled questions from the test set is around \$0.34.
@@
A.6 FuSIC-KBQA Details
Our proposed approach FUn-FuSIC builds upon the the base architecture of FuSIC-KBQA. FuSIC-KBQA has a three step pipeline: (a) Supervised Retrieval: a supervised retriever, trained on the source domain and optionally fine-tuned on the target domain is used to obtain the top-100 classes, relations and paths that are relevant to the question asked, (c) LLM Generation: We provide the top-10 classes, top-10 relations and top-5 paths along with few-shot exemplars to generate the sparql query.
A.7 FUn-FuSIC Prompts
@@
A.8 PUn prompt
The first prompt is for Prompting for Unanswerability (PUn).
@@ A.8 PUn prompt
A.8.1 FUN prompt
@@
The next three prompts fall under the answer incompatibility feedback
@@
A.8.2 Prompt for Verifier V3 Question Logical Form Agreement
The few shots provided for verifying question logical form agreement come from the few shots provided in the target domain $D^{t}$. We obtain positive samples from the dataset $D^{t}$ directly, using the questions and gold logical forms. For obtaining negative samples, we perform zero-shot FuSIC-KBQA inference over the dataset $D^{t}$. Then we consider those questions for which the predicted logical form is different from the gold logical form.
Firstly, we perform back-translation to obtain natural language question from the logical form
@@ A.8.2 Prompt for Verifier V3 Question Logical Form Agreement
We use few-shot LLM prompting to obtain the explanation for why the question and logical form agree or disagree. These few-shots, for obtaining the explanation are dataset independent, and are manually written.
@@
Some examples of few shots are provided below-
Question Logical Form Sgreement Check(V3(iii)) Question we answer: Who are the cricket players who have made exactly 31 stumps in one day internationals? Question originally asked: name the cricket player who has 31 odi stumps. explanation: The question we answer returns cricket players. The question originally asked also returns cricket players. The question we answer finds cricket players who have made exactly 31 stumps in one day internationals. The question originally asked also finds cricket players who have made 31 stumps in one day internationals. Both questions involve no mathematical or logical operators. Hence, they are same. Question we answer: Which game engines are successors to the Unreal Engine? Question originally asked: which video game engine’s successor engine is unreal engine? explanation: The question we answer returns game engines. The question originally asked also returns game engines. The question we answer finds successors to the Unreal Engine. The question originally asked finds the predecessor of the Unreal Engine. The reasoning steps followed by the two questions are different. Hence, they are different. [total 6 exemplars] Question we answer: Which cars drive at a speed of 80? Question originally asked: name the car with driving speed at least 80? explanation:
A.8.3 Prompt in scUn for Logical Form Selection
We store the back-translated natural language questions along with each predicted logical form. Here, we prompt the LLM to select the closest back-translated natural language question to the original question.
@@ A.8.3 Prompt in scUn for Logical Form Selection
A.9 Supervised Models training details
We use Hugging Face  Wolf et al. (2020), PyTorch Paszke et al. (2019) for our experiments and use the Freebase setup specified on github 777 https://github.com/dki-lab/Freebase-Setup . We use NVIDIA A100 GPU with 40 GB GPU memory and 32 GB RAM. For training the discriminator module of RetinaQA, we require 2 GPUs. (1) For the answerable experiments, we use the supervised models as specified in Patidar et al. (2024). (2) For the unanswerability experiments, we train all models from scratch. (a) We use RnG-KBQA entity linker  [^8] https://github.com/salesforce/rng-kbqa/tree/main/GrailQA/entity_linker (BSD 3-Clause License) trained on the answerable subset of GrailQAbility for all our experiments. (b) We train the RnG-KBQA path retriever on answerable subset of WebQSP [^9] https://github.com/salesforce/rng-kbqa/blob/main/WebQSP/scripts/run_ranker.sh (BSD 3-Clause License). The number of training epochs is determined by the performance of the model over the answerable questions in the dev set. (c) We train the TIARA schema retriever on the answerable subset of WebQSP  [^10] https://github.com/microsoft/KC/tree/main/papers/TIARA/src (MIT License) (d) We train the sketch generator and discriminator of RetinaQA on the answerable subset of WebQSP [^11] https://github.com/dair-iitd/RetinaQA.
A.10 Supervised Models inference details
We train all components on WebQSP, using the corresponding target domain’s dev set as a validation set for early stopping. In the absence of unanswerable questions for training, both models use a threshold fine-tuned on a dev set to detect schema-level unanswerability. We again use the target dev sets for this.
We use the dev set in RetinaQA, during discriminator inference for (a) determining how to best utilize the candidate paths. We might (i) not provide candidate paths (ii) provide candidate paths in GrailQA format (iii) provide candidate paths in WebQSP format. We select the best alternative based upon the performance of the model over the dev set. For the WebQSP $\rightarrow$ GrailQAbility dataset, we observe (ii) works best, whereas for the WebQSP $\rightarrow$ GraphQAbility dataset, we observe (i) works best. (b) determining the threshold value. RetinaQA applies a threshold on the scores - for a question, if the highest score candidate logical form has a score less than the threshold, the question is labeled as NK. We choose the optimal value of the threshold to maximize the overall EM-s score over the dev set.
A.11 Pangu Adaptation Details
Similar to RetinaQA, We train all components on WebQSP, using the corresponding target domain’s dev set as a validation set for early stopping. We use 1 GPU for training. Same as RetinaQA, we use the dev set to determine the threshold for schema-level unanswerability. Pangu-T applies a threshold on the scores - for a question, if the highest score candidate logical form has a score less than the threshold, the question is labeled as NK. We choose the optimal value of the threshold to maximize the overall EM-s score over the dev set.
A.12 KB-Binder and FuSIC-KBQA Adaptation Details
For KB-Binder, we make use of publicly available code  [^12] https://github.com/ltl3A87/KB-BINDER (MIT License). We use self-consistency and majority voting with 6 examples, as in the experiments in the paper. In the retrieval(-R) setting, KB-Binder samples demonstration examples by retrieving from the entire available training data. We restrict its retrieval to our target training set $D_{t}$ with 25 examples. KB-Binder reports experiments using code-davinci-002 as the LLM. For consistency and fair comparison, we replace these with gpt-4-0613 as in our approach. This model generates logical forms in s-expression, which we preserve.
For FuSIC-KBQA Patidar et al. (2024), since no code is available for this model, we use our own implementation based on the description in the paper. For FuSIC-KBQA, and FUn-FuSIC we use $temperature=0$.
A.13 Real Example of FUn
Question: Where all were the Olympics held before 2024?
Logical Form 1 ?y{ ?x olympics.venue ?y
    ?y olympics.year ?z
    FILTER(?z = 2024)}
Feedback 1 Type: Logical Form Semantic Error Description: ?y is assigned types [’venue’, ’olympics’] by relations [’olympics.venue’, ’olympics.year’]. These types are mutually incompatible.
Logical Form 2 ?y{ ?x olympics.venue ?y
    ?x olympics.year ?z
    FILTER(?z = 2024)}
Feedback 2 Type: Natural Language Semantic Error Description: The question originally asked returns all places where the Olympics were held before 2024. However, the question we answer returns the places where the Olympics were held in 2024.
Logical Form 3 ?y{ ?x olympics.venue ?y
    ?x olympics.year ?z
    FILTER(?z < 2024)}

### v3

@@ Appendix A Appendix
A.1 KBQA Elaboration
Here we elaborate on different aspects of the KBQA problem.
A.1.1 Challenges in Few Shot Transfer Learning for KBQA
The source and target tasks may differ significantly. First, the data and schema of the knowledge bases $G^{t}$ and $G^{s}$ and the domains they cover may be different. Secondly, the distributions of questions and logical forms defined over the KBs may be different in $D^{t}$ and $D^{s}$.
A.1.2 Different Types of Unanswerability
@@
A.2 Algorithm for Self Consistency with Unanswerability
[…]
The high level algorithm for assessing confidence in the set of candidate logical forms (assessConf) is described in Algo. 4. We use abbreviations NE and E to indicate non-empty and empty respectively. popAnsNE is abbreviation of ”(most) popular answer non-empty”.
Algorithm 4 $\mbox{assessConf}(q,L,\mathcal{L})$
[…]
2:if $(c)$ then
3:  $l=\mbox{selectBestNE}(q,L^{p},\mathcal{L})$
4:  $\mbox{{\bf return}}(\mbox{T},l,A^{p})$
5:end if
[…]
7:if $(c)$ then
8:  $l=\mbox{selectBestE}(q,L^{p},\mathcal{L})$
9:  $\mbox{{\bf return}}(\mbox{T},l,\texttt{NA})$
10:end if
[…]
A.3 Details of FUn Verifiers
Here we discuss the verifiers V2, V3 and V4 in more detail.
@@ A.3 Details of FUn Verifiers
(V3) Question-Logical form Disagreement:
FUn performs equivalence check between $l$ and $q$ using a novel multi-stage LLM pipeline. (i) The variable names in $l$ are first naturalized to $l^{n}$ considering $q$ and preserving semantics, e.g. by replacing ’?x’ with ’?actor’. (ii) $l^{n}$ is back-translated into a natural language question $q^{b}$. (iii) $q^{b}$ is finally checked for semantic equivalence with $q$. The first two steps are performed using zero-shot prompting, while the last is performed using few-shots constructed using the target few-shots $D^{t}$. The feedback mentions lack of equivalence as the type of error.
@@
Knowledge Base Logical Forms NL Questions
[…]
SrcAv/TgtAv Cosine Sim SrcAv/TgtAv
[…]
Table 6: Statistics for different source and target (test set) KBQA task pairs in terms of the knowledge base, logical forms and natural language questions. ‘Domain(JS)’ is Jensen Shannon (JS) divergence between domain distribution of questions, ‘New Rel%’ is percentage of questions in target with new (unseen) relations, ‘Function (JS)’ is JS-divergence between distributions over functions in logical forms, ‘#\##Relations’ shows the source average and the target average for number of relations per logical form, ‘Src-Tgt Cosine Sim’ is average minimum cosine distance between source and target questions and ‘#\##Tokens’ shows the source average and target average of number of tokens per question.
A.4 Additional Dataset Statistics
[…]
Dataset Schema level Data level
UnAns UnAns
[…]
Table 7: Percentages of schema and data level unanswerability among unanswerable questions.
A.5 EM-s: Automated Approximate Equivalence Check for SPARQL
@@
As has been observed in Patidar et al. (2023), answer evaluation by itself is not a robust measure for evaluation of KBQA models when the dataset contains unanswerability. Traditional KBQA models that generate s-expressions can be evaluated using EM, which checks for logical form equivalence between two logical forms, since it is possible to compare equivalence between two s-expressions efficiently. However, FUn-FuSIC generates SPARQL queries instead. Directly comparing program equivalence between two SPARQL queries is an undecidable problem  [^7] https://users.dcc.uchile.cl/~cgutierr/papers/expPowSPARQL.pdf. Patidar et al. (2024) suggest a semi-automatic strategy for comparison of sparql queries. We propose a completely automatic metric for SPARQL equivalence check. Two SPARQL queries are equivalent by the EM-s check if (a) the relations occurring in the two queries are same. (b) the entities occurring in the two queries are the same (c) the answer set obtained by executing the queries over the KB are the same. Note that the EM-s check is necessary, but not sufficient for two SPARQL queries to be equivalent.
[…]
A.6 Model Evaluation: Additional Details
@@
In this section, we do a deeper evaluation of performance of different models across different categories of unanswerability, as explained in Patidar et al. (2023). There are two broad categories of unanswerability — schema level unanswerability (absence of knowledge in terms of KB ontology or entities required to construct the logical form) and data level unanswerability (absence of facts or intermediate entities of the logical form path on the KB).
@@
We expect that (a) due to poor ability of supervised models to generalize in transfer learning settings, RetinaQA will struggle to generate correct logical forms for data level unanswerable questions, and (b) due to the strong generalization ability of FuSIC-KBQA, it should be able to perform well for data level unanswerable questions. However, since it is biased towards returning incorrect logical forms instead of abstaining from returning a logical form, it will perform poorly at identifying schema level unanswerable questions. (c) FUn-FuSIC should be able to maintain the performance of FuSIC-KBQA on data level unanswerable questions to a large extent, while significantly improving the performance on schema level unanswerable questions.
[…]
A.7 FUn Cost Analysis for Proprietary LLMs
@@
FuSIC-KBQA, as well as the adapted versions of FuSIC-KBQA, such as FuSIC-KBQA-U and FUn-FuSIC rerank the classes, relations and paths. The total cost for reranking for one question is \$0.16. The cost for generation of logical form from a prompt with 5 in-context examples is \$0.16. Thus, the approximate cost for inference of one question by FuSIC-KBQA-U is \$0.32.
@@
The cost for generation of logical form from a prompt with 0 in-context examples is \$0.04. The cost of checking whether two natural language questions are equivalent or not, using few-shot exemplars and chain of thought prompting is also \$0.04. The approximate cost of inference of one question by FUn-FuSIC varies between \$0.24 and \$0.48. The average cost over 50 randomly sampled questions from the test set is around \$0.34.
@@
A.8 Model Adaptation Details
Here we discuss adaptation details for the models that we have built upon (FuSIC-KBQA), used as retrievers (RetinaQA) and for comparison.
A.8.1 FuSIC-KBQA Details
Our proposed approach FUn-FuSIC builds upon the the base architecture of FuSIC-KBQA Patidar et al. (2024). FuSIC-KBQA has a three step pipeline: (a) Supervised Retrieval: a supervised retriever, trained on the source domain and optionally fine-tuned on the target domain is used to obtain the top-100 classes, relations and paths that are relevant to the question asked, (c) LLM Generation: We provide the top-10 classes, top-10 relations and top-5 paths along with few-shot exemplars to generate the SPARQL query.
Since no code is available for this model, we use our own implementation based on the description in the paper. For FuSIC-KBQA, and FUn-FuSIC we use LLM $temperature=0$.
A.8.2 Training Details for Supervised Models
We use Hugging Face  Wolf et al. (2020), PyTorch Paszke et al. (2019) for our experiments and use the Freebase setup specified on github  [^8] https://github.com/dki-lab/Freebase-Setup . We use NVIDIA A100 GPU with 40 GB GPU memory and 32 GB RAM. For training the discriminator module of RetinaQA, we require 2 GPUs. (1) For the answerable experiments, we use the supervised models as specified in Patidar et al. (2024). (2) For the unanswerability experiments, we train all models from scratch. (a) We use RnG-KBQA entity linker  [^9] https://github.com/salesforce/rng-kbqa/tree/main/GrailQA/entity_linker (BSD 3-Clause License) trained on the answerable subset of GrailQAbility for all our experiments. (b) We train the RnG-KBQA path retriever on answerable subset of WebQSP [^10] https://github.com/salesforce/rng-kbqa/blob/main/WebQSP/scripts/run_ranker.sh (BSD 3-Clause License). The number of training epochs is determined by the performance of the model over the answerable questions in the dev set. (c) We train the TIARA schema retriever on the answerable subset of WebQSP  [^11] https://github.com/microsoft/KC/tree/main/papers/TIARA/src (MIT License) (d) We train the sketch generator and discriminator of RetinaQA on the answerable subset of WebQSP [^12] https://github.com/dair-iitd/RetinaQA.
A.8.3 Inference Details for Supervised Models
We train all RetinaQA components on the source WebQSP’s training set, using the corresponding target domain’s dev set as a validation set for early stopping. In the absence of unanswerable questions for training, both models use a threshold fine-tuned on a dev set to detect schema-level unanswerability. We again use the target dev sets for this.
We use the dev set in RetinaQA, during discriminator inference for different purposes. (A) Determining how to best utilize the candidate paths. The possibilities are (i) not providing candidate paths, (ii) providing candidate paths in GrailQA format, and (iii) providing candidate paths in WebQSP format. We select the best alternative based upon the performance of the model over the dev set. For the WebQSP $\rightarrow$ GrailQAbility dataset, we observe (ii) works best, whereas for the WebQSP $\rightarrow$ GraphQAbility dataset, we observe (i) works best. (B) Determining the threshold value. RetinaQA applies a threshold on the scores - for a question, if the highest score candidate logical form has a score less than the threshold, the question is labeled as NK. We choose the optimal value of the threshold to maximize the overall EM-s score over the dev set.
A.9 Pangu Adaptation Details
Similar to RetinaQA, we train all Pangu components on WebQSP, using the corresponding target domain’s dev set as a validation set for early stopping. We use one GPU for training. Same as RetinaQA, we use the dev set to determine the threshold for schema-level unanswerability. Pangu-T applies a threshold on the scores - for a question, if the highest score candidate logical form has a score less than the threshold, the question is labeled as NK. We choose the optimal value of the threshold to maximize the overall EM-s score over the dev set.
A.9.1 KB-Binder Adaptation Details
For KB-Binder, we make use of publicly available code  [^13] https://github.com/ltl3A87/KB-BINDER (MIT License). We use self-consistency and majority voting with 6 examples, as in the experiments in the paper. In the retrieval(-R) setting, KB-Binder samples demonstration examples by retrieving from the entire available training data. We restrict its retrieval to our target training set $D_{t}$ with 25 examples. KB-Binder reports experiments using code-davinci-002 as the LLM. For consistency and fair comparison, we replace this with gpt-4-0613 as in other LLM-equipped models. KB-Binder generates logical forms in s-expression, which we preserve.
WebQSP $\rightarrow$ WebQSP $\rightarrow$
Model GrailQAbility GraphQAbility
V3(a) 92 96
V3(b) 94 90
V3(c) 92 94
V3(overall) 88 90
V4(b) 75 68
Table 8: Accuracy of weak verifier on the two datasets by manual analysis of 100 instances from the test sets.
A.10 Accuracy of Weak Verifiers
Accuracy of the two weak verifiers are recorded in detail in Tab. 8.
A.11 FUn-FuSIC Prompts
@@
A.11.1 PUn prompt
The following prompt is for Prompting for Unanswerability (PUn).
@@ A.11.1 PUn prompt
A.11.2 FUN Prompts
The following is the prompt used by FUn for accommodating feedback from verifier V1.
@@ A.11.2 FUN Prompts
The following is the prompt used by FUn for accommodating feedback from verifier V2.
@@
The following is the prompt used by FUn for accommodating feedback from verifier V3.
@@
The following three prompts are used by FUn for accommodating feedback from verifier V4.
@@
A.11.3 Prompt for Question Logical Form Agreement Verifier (V3)
The few shots provided for verifying question logical form agreement are derived from $D^{t}$. We obtain positive samples from the dataset $D^{t}$ directly, using the questions and gold logical forms. For obtaining negative samples, we perform zero-shot FuSIC-KBQA inference over $D^{t}$. Then we consider those questions for which the predicted logical form is different from the gold logical form.
First, we perform back-translation to obtain natural language question from the logical form using the following prompt.
@@ A.11.3 Prompt for Question Logical Form Agreement Verifier (V3)
We use few-shot LLM prompting to obtain the explanation for question and logical form agreement or disagreement.
@@
These few-shots for obtaining the explanation are dataset independent, and are manually written. Some examples of few shots are below.
Question Logical Form Agreement Check(V3(iii)) Question we answer: Who are the cricket players who have made exactly 31 stumps in one day internationals? Question originally asked: name the cricket player who has 31 odi stumps. explanation: The question we answer returns cricket players. The question originally asked also returns cricket players. The question we answer finds cricket players who have made exactly 31 stumps in one day internationals. The question originally asked also finds cricket players who have made 31 stumps in one day internationals. Both questions involve no mathematical or logical operators. Hence, they are same. Question we answer: Which game engines are successors to the Unreal Engine? Question originally asked: which video game engine’s successor engine is unreal engine? explanation: The question we answer returns game engines. The question originally asked also returns game engines. The question we answer finds successors to the Unreal Engine. The question originally asked finds the predecessor of the Unreal Engine. The reasoning steps followed by the two questions are different. Hence, they are different. [total 6 exemplars] Question we answer: Which cars drive at a speed of 80? Question originally asked: name the car with driving speed at least 80? explanation:
A.11.4 scUn Prompt for Logical Form Selection
We store the back-translated natural language questions along with each predicted logical form. We use the following prompt to select the closest back-translated natural language question to the original question.
@@
A.12 Example of FUn Iterations
The following shows 3 Fun iterations for an example question from the test set of WebQSP $\rightarrow$ GrailQAbility dataset.
Question: what is the musical genre of the recording who m i (feat. 일리닛, new champ, myk)?
Entities: who m i (feat. 일리닛, new champ, myk) m.0123lk0s
Logical Form 1 ?x {
    :m.0123lk0s :music.genre.recordings ?x
    ?x :type.object.type :music.genre}
Feedback 1 Type: Logical Form Semantic Error Description: The types of relations don’t match for entity in the query. The assigned relation types by [’music.genre.recordings’] are [’music.genre’]. These types are not associated with this entity in the KB.
Logical Form 2 ?y {
    :m.0123lk0s :music.recording.artist ?x
    ?x :music.artist.genre ?y
    ?y :type.object.type :music.genre}
Feedback 2 Type: Natural Language Semantic Error Description: The question originally asked genre of the song. However, the question we answer returns genre associated with artist of the song.
Logical Form 3 ?x { ?x
    :music.genre.recordings :m.0123lk0s
    ?x :type.object.type :music.genre}